OpenStudy (anonymous):

What is the length of side BC? (3,1 to 1,1)

OpenStudy (anonymous):

\[d = \sqrt{(x_{1} - x_{2})^{2} + (y_{1} - y_{2})^{2}}\]\[d = \sqrt{(3 - 1)^{2} + (1 - 1)^{2}}\]Can you try from here?

OpenStudy (anonymous):

iI have no idea what time is is where your at, but its like: 9:33 pm here, and i really need answers...i feel bad, and i really need answers i've tried everything even calulators for this...

OpenStudy (anonymous):

Do what's in the parentheses first. 3 - 1 = ?

OpenStudy (anonymous):

two.

OpenStudy (anonymous):

4+0?

OpenStudy (anonymous):

Ok. 1 - 1 = 0 \[d = \sqrt{2^2} = 2\]

OpenStudy (anonymous):

Yup. \[d = \sqrt{4 + 0}\]

OpenStudy (anonymous):

yay -.- 5 more questions to go.lucky me. -___-

OpenStudy (anonymous):

lol. Private message me the questions. I'll help you lickity split. (did I just say lickity split?)

OpenStudy (anonymous):

& i can't move on to see them all, untill i answer then BLEHEHEHEHEH